OpenStudy (anonymous):

the product of three consecutives integers n-1,n,and n+1 is -336.what are the integers?HELP PLEASE

OpenStudy (anonymous):

n-1 + n + n+1 = 3n 3n=-336 n=-113

OpenStudy (anonymous):

THANK YOU!!!

OpenStudy (anonymous):

(n - 1)(n)(n + 1) = -336 which gives: n^3 - n + 336 = 0 Synthetic division gives n = -7, so your numbers are -6, -7, and -8

OpenStudy (anonymous):

Aha, sorry ignore my answer @tcarroll010 has the right one, I thought it was a sum and not a product

OpenStudy (anonymous):

can u show the step in synthetic division please

OpenStudy (anonymous):

np @SmartWish you were definitely on the right track and just had a little lapse.

OpenStudy (anonymous):

no problem,thank you for help me in this anyway!!!

OpenStudy (anonymous):

You can go through synthetic division, or a much simpler intuitve way is to get an approximation of taking the third root of -336 and you will see that it is very close to -7.

OpenStudy (anonymous):

yes you are definetily right

OpenStudy (anonymous):

You could go through synthetic division to get the answer, but, believe it or not, you basically have to know the answer before actually doing the synthetic division! So, in a case like this, it is much much much better to do the approximation.

OpenStudy (anonymous):

thanks you help me a lot

OpenStudy (anonymous):

Good luck to you and thx for the recognition!

OpenStudy (anonymous):

ur welcome!1!!